Operating Cash Flow

The financial staff of Cairn Communications has identified the following information for the first year of the roll-out of its new proposed service:

Projected sales $20 million
Operating costs (not including depreciation) 7 million
Depreciation 4 million
Interest expense 3 million

The company faces a 35% tax rate. What is the project's operating cash flow for the first year (t = 1)? Write out your answer completely. For example, 2 million should be entered as 2,000,000.
Answer #1
Calculations Particulars Amount $
A
Projected sales
20000000
B Operating costs (not including depreciation) 7000000
C Depreciation 4000000
D = A-B-C Profit Before Interest and Taxes 9000000
E Interest expense 3000000
F = D - E Profit Before Taxes 6000000
G = F x 35% Taxes 2100000
H = F - G Profit After Taxes 3900000
I = C + H Operating Cash Flow 7900000

> This is incorrect

Frank Abagnale Sun, Dec 5, 2021 7:05 PM
Answer #2

(Projected Sales - Operating costs (not including depreciation))*(1-Tax Rate)+(Depreciation*Tax Rate)        

(20,000,000-7,000,000)*.65+(4,000,000*0.35)        

=9,850,000

answered by: Frank Abagnale